OpenStudy (anonymous):

how to do you look for the y-coordinate or 'b' in the slope-intercept form equation: y=mx+b?

OpenStudy (imer):

"m" is the slope of the line "b" is the y-intercept The equation of form "y=mx+b" is a linear equation

OpenStudy (anonymous):

the question is "A line passes through (0,0) and (1,1). Find the slope." I dont know what to use and how to solve this thing

OpenStudy (anonymous):

m=y2-y1/x2-x1

OpenStudy (imer):

The slope is the measure of steepness of a line. And we measure that by finding the ratio of difference in "y-coordinate" and "x-coordinate"

OpenStudy (imer):

Correct, you have been given two points, use the formula you gave me to find the slope.

OpenStudy (anonymous):

my answer is 1/1 and simplified it to 1. so m=1 is my answer correct?

OpenStudy (imer):

Correct!, so what should be the equation of a line that passes through point (0, 0) and (1, 1)?

OpenStudy (anonymous):

am i supposed to use y=mx+b? (just guessing, im sorry)

OpenStudy (imer):

yes, you have to replace "something" in y=mx+b, what letter will you replace with the slope you found?

OpenStudy (anonymous):

m :D so it'll be y=1x+b? :D

OpenStudy (imer):

correct, but now what is the y-intercept which is "b"?

OpenStudy (anonymous):

that's the problem :( i dont know how to look for the y-intercept :(

OpenStudy (imer):



OpenStudy (imer):

And when a line passes the y-axis, what is x-coordinate at the point?

OpenStudy (anonymous):

so the y-intercept is 1? (lol i apologize for my stupidity)

OpenStudy (imer):



OpenStudy (imer):

we know when the line cuts y-axis, x-coordinate is "0" But that was a mere example. Tell me from the points you have been given, what is y-intercept?

OpenStudy (anonymous):

1? lol

OpenStudy (imer):

You are given two point (0, 0) and (1, 1) , in which point x-coordinate is "0"?

OpenStudy (anonymous):

oh. 0 then lol

OpenStudy (anonymous):

so it'll be y=1x+0??

OpenStudy (imer):

Correct! because x-coordinate is "0" in the first point, therefore, you look for y-coordinate in that point but y-coordinate is "0" too, therefore, y-intercept is "0"

OpenStudy (imer):

correct, and if you simplify you get \[y=1x+0 \rightarrow y=x\]

OpenStudy (imer):

and then if you feel you made a mistake you can verify if the line passes through the points you are given, when x=0, y=0 and when x=1, y=1 therefore the equation is correct.

OpenStudy (anonymous):

Ah! is y=1x-0 the final answer or there's more? its not in the choices tho :(

OpenStudy (imer):

Yes, it's the final answer in form y=mx+c

OpenStudy (anonymous):

but its not in the choices :(

OpenStudy (imer):

if the question is asking you to find the slope, then the answer is "1", and if the question is asking you to find the y-intercept then the answer is "0"

OpenStudy (imer):

the question is "A line passes through (0,0) and (1,1). Find the slope." Therefore, the answer is "1"

OpenStudy (imer):

You mentioned "y=mx+c" in your original question so I thought you need the equation of the line.

OpenStudy (anonymous):

Sorry Late reply. THANK YOU SO MUCH!! YOU ROCK :))